OpenStudy (anonymous):

At the 2006 olympics Germany won a total of 29 medals they won 1 more silver medal than gold and 5 more gold medals than bronze. Find the number of each type of medals?

OpenStudy (paounn):

You have three variables, (the number of gold, silver, and bronze medals, g, s, b rispectively) and three conditions: 1: sum of all medals is 29: g+s+b= 39 2: 1 more silver than gold: s+1= g 2: 5 more gold than bronze: b+5= g

OpenStudy (anonymous):

2 should be s = g + 1 because which is greater, # of silver or # of gold?

OpenStudy (anonymous):

oh, and "respectively"

OpenStudy (anonymous):

Solve the following for g, s and b, the number of gold, silver and bronze medals respectively.\[\{b+g+s=29,s=1+g,g-5=b\}\]\[\{b\to 6,g\to 11,s\to 12\} \]

OpenStudy (anonymous):

thank you robtobey

OpenStudy (anonymous):

You're welcome.